# Parallel Program Analysis on Path Ranges

### v1

@@ 4 Joining Witnesses
First, let us look at the verdict “false”, i.e., the ranged program analysis reports that the program violates the analyzed property. For the verdict “false”, it is sufficient that at least one of the ranged analyses reports “false”. This analysis also produces a violation witness [18, 19] that describes a counterexample path, which witnesses the violation. Furthermore, its violation witness is sufficient to explain the aggregated verdict “false” of the ranged program analysis. Moreover, providing more than one counterexample path does not make for a better explanation of the aggregated verdict. Hence, we decided against joining violation witnesses and simply use one of the violation witnesses provided by a ranged analysis with result “false”. More concretely, we use the first violation witness being generated.
@@
Next, let us consider the verdict “true”, i.e., the ranged program analysis proves the property. In this case, all ranged analysis successfully proved that their subset of the program behavior fulfills the property. To explain their results, they each provide a correctness witness [11, 19]. Formally, a correctness witness for a CFA is a (protocol) automaton . Similar to most automata, a correctness witness consists of a set of states including initial state , an alphabet whose symbols define which operations (i.e., CFA edges) and data states are considered by an automaton transition, and a transition relation . In addition, it contains a total function that assigns to each state of the automaton an invariant describing an expected set of data states. Furthermore, a correctness witness must fulfill the following completeness property, which guarantees that the correctness witness does not exclude any program behavior.
@@
Figure 4 shows three correctness witnesses for our example program. For this example, their graph structure is identical and reflects the CFA structure. However, the witnesses differ in their invariants. The invariants are the second part of the node labels and are shown in a different color for each witness. Note that we use o/w (short for otherwise) to summarize all control-flow edges that do not occur in another outgoing edge of the respective state. For example, for the initial state o/w stands for . The witnesses with the green and red invariants, respectively, both provide an explanation for the correctness of the subset of the program behavior. The witness with green invariants might be generated by a ranged analysis based on predicate abstraction [20] considering range [] resulting from test . The witness with the red invariants might be produced by a ranged analysis based on value analysis [21] considering the opposite range []. We observe that both witnesses use invariant “false” for locations that the respective analysis never explored because no paths in their range contain these locations. Furthermore, note that the analyses use block encoding and, therefore, may only provide invariants for particular locations (either loop heads or locations with multiple outgoing edges). Locations, which are intermediate locations in the block encoding, are assigned the invariant true.
@@
Each of the two witnesses generated by a ranged analysis provides an explanation for the correctness of the subset of the program behavior analyzed by the respective ranged analysis but likely does not provide sufficient information for behavior analyzed by another ranged analyses. In the following, we explain how to combine correctness witness produced by ranged analyses during ranged program analysis into a joint correctness witness that explains the correctness of the complete program. The black invariants belong to such a joint witness. In the following, we explain how to compute these joint witnesses.
@@
Algorithm 1 describes the combination for two correctness witnesses888The algorithm can easily be extended to combine more than two witnesses. Alternatively, one can for example iteratively combine multiple witnesses by first joining the first two witnesses with Alg. 1 and then successively using Alg. 1 to combine the result of the first n-1 witness combinations with the n-th witness.. Its main part (lines 1–9) performs a parallel exploration of the program and the two witnesses. While a parallel exploration of the two witnesses would have been sufficient to compute a joint witness, we include the program into the exploration because it allows us to focus on explanations for program behavior, to construct (compact) joint witnesses that align with the program structure, which we assume are better for validation, and to simplify the witness construction. The parallel exploration starts in the initial location and initial automata states and explores the successors of each explored node once. Since we only need to provide explanations for program behavior, i.e., program paths, successor exploration focuses on successors that adhere to the program’s control-flow. To this end for each leaving control-flow edge in the program, we inspect all combinations of outgoing transitions of the witnesses that consider edge and that can be taken jointly, i.e., their enabling conditions do not exclude each other. The resulting successor combinations are then considered further during exploration. After the parallel exploration, the joint witness is computed. To avoid that the joint witness forces a validator to unroll the program, which may make it inefficient, Alg. 1 computes a rather natural representation of the joint correctness witness, which aligns with the program structure. Therefore, line 10 computes one joint invariant per program location. For each node detected during the parallel exploration, we combine the invariant information from both witnesses. We choose to combine the information via disjunction. Thus, we may loose information in case both witnesses considered the full behavior when computing their invariant information. However, disjunction ensures that we remain sound in case different behaviors are considered. Based on the combination, the joint invariant for a program location is derived from the disjunction of all nodes explored that refer to that location . We use a disjunction because different nodes likely represent different behavior. Note that we include “false” to the disjunction to account for syntactically unreachable locations. After computing the invariants, line 11 determines the transitions of the joint witness. We add transitions to encode the program’s control-flow. Due to the format of a transition, we cannot add the control-flow edge directly, but add a transition that accepts the control-flow edge and that is always enabled. To make the witness complete, we add self-loops to each state999More precisely, we add self-loops to each program location of the joint witness that cover all remaining edges, which do not adhere to the control-flow. Lastly, line 12 returns the joint witness, which reuses the program’s locations and initial location as states and initial states, uses the same alphabet as the witnesses and the previously computed transitions and joint invariants.
@@
We implemented the Alg. 1 for joining of witnesses presented above in the software analysis framework CPAchecker [12]. For the parallel exploration, we compose a CPA tracking the control-flow with one CPA per witness that tracks the transitions of that witness. To allow us to combine the witness information of the explored states, we wrap the above composition of CPAs into another CPA (called WrapperCPA), whose transfer relation forwards to the composition. We then use CPAchecker’s meta-reachability analysis instantiated with the above WrapperCPA to perform the parallel exploration. Thereafter, we apply CPAchecker’s witness generation on the result of the parallel exploration, which relies on the WrapperCPA to combine the witness information and computes the witness similar to Alg. 1.
@@
Finally, let us briefly discuss which invariants a ranged analysis should provide for locations it never explored. Basically, there are two options: “true” or “false”. In contrast to “false”, “true” is always a correct invariant with respect to the complete program. However, correctness witnesses of ranged analyses that only analyze a subset of the program behavior already do not guarantee correct invariants for the complete program for locations they visited during their analysis, especially if they do not analyze all paths to that location. They only encode information on behavior that they have seen. Thus, “false” would better fit to the information encoded at other seen locations. In addition, providing “true” would result in loosing all correctness witness information provided by a different ranged analysis due to the disjunction. Hence, from the perspective of witness joining “false” is preferred. Since correctness witnesses were not designed for incomplete analysis and do not provide any requirements on how to handle unreached code, it depends on the verifier how it handles locations not analyzed. Hence, the quality and usefulness of a joint correctness witness depends on how the individual verifiers in the ranged program analysis handle non-analyzed locations and, therefore, may differ amongst verifiers.
@@ 6 Implementation
To show the advantages of ranged program analysis and to evaluate the two novel techniques of witness joining and work stealing, we realized ranged program analysis and the two novel extensions within CoVeriTeam [23], a tool for specifying compositions of verification tools.
6.1 Realization of Ranged Program Analysis
At first, we realized the composition of range reduction from Sec. 3.2 with an arbitrary configurable program analysis in the tool CPAchecker [12], making use of the existing composite pattern. The realization of the range reduction follows our formalization, i.e., it reuses elements from the value analysis, which are already implemented within CPAchecker. For the use of non-CPA-based analyses, we refer to [24], where a ranged program analysis based on instrumentation is presented, that allows for using any off-the-shelf analysis, making use of reduction [3] and program instrumentation.
@@ 6.1 Realization of Ranged Program Analysis
The composition of ranged program analyses from Sec. 3 is realized using the tool CoVeriTeam [23]. The tool allows building parallel and sequential compositions making use of existing program analyses, especially those from the CPAchecker. We use CoVeriTeam for the orchestration of the composition of ranged analyses, especially for work stealing and the aggregation of the partial results. To be able to realize work stealing, i.e. to keep track of which program range is already successfully analyzed, we build a novel ranged program analysis component within CoVeriTeam. It follows the description in Fig. 2 and Sec. 5 and extracts the verdicts and witnesses generated. Beneath the two ranged analyses Ranged Analysis 1 and Ranged Analysis 2, the implementation takes a splitter as input. We realized the splitters Lb3 described in Sec. 3.4 as standalone components within CPAchecker. It generates test cases in the standardized XML-based TEST-Comp test-case format101010https://gitlab.com/sosy-lab/test-comp/test-format/blob/testcomp23/doc/Format.md. If the splitter fails, e.g. Lb3 cannot compute a test case, when the program does not contain a loop, we execute Ranged Analysis 1 on the interval . The implementation of ranged program analysis can be configured to join the partial verification witnesses generated by Ranged Analysis 1 and Ranged Analysis 2. For this, we implemented Alg. 1 as a standalone component within CPAchecker.
6.2 Employed CPA-based Analysis.
Value Analysis. Value Analysis [21] (sometimes called constant propagation or explicit state model checking) tries to compute the concrete values for variables occurring in the program. For this, the exact value of each variable from a set (called precision) is tracked along each program path. Information for locations reachable via multiple program paths are not merged. The employed configuration does not track all variable values but uses counterexample-guided abstraction refinement (CEGAR) [25] to iteratively refine the precision until the program can either be proven safe or a violation is detected.
Predicate Abstraction. The predicate abstraction used is the standard one in the CPAchecker, configured to compute an abstraction of the program using predicates with adjustable block encoding [20], abstracting only at loop heads. The predicates used for the analysis are determined by CEGAR [25], lazy refinement [27], and interpolation [28].
Bounded Model Checking. The configuration of bounded model checking (BMC) employed is iterative, meaning that the -th iteration inspects the behavior of the CFA by unrolling the loop to a bound of iterations. If no property violation can be found, is increased. For inspecting the behavior of the program, BMC encodes the unrolled CFA and the property as a boolean formula using the unified SMT-based approach for software verification [29]. Hereafter, the satisfiability of the formula that encodes the program up to the bound is checked to detect property violations. In case that is large enough that the full program is encoded, and no property violation is detected, the program is safe.
@@ 7 Evaluation
The evaluation of ranged symbolic execution conducted by Siddiqui and Khurshid [10] focuses on the performance increases gained for symbolic execution. We showed in [14] that ranged program analysis can increase the performance of a single analysis (for symbolic execution and partially for BMC) and that the combination of two analyses (symbolic execution and predicate abstraction) increases the performance for symbolic execution, the one with the lower performance. In addition, we compared the performance of different splitters, whereas Lb3 leads to the best overall performance. In the following, we are interested in a more systematic analysis of the effect of using ranged program analysis. We first want to investigate, if using two instances of the same analysis within ranged program analysis pays off for all four basic analyses. Next, we study the effect of combining two conceptually different analyses and also analyze work stealing. Lastly, we study the question of whether our ranged program analysis is still capable of generating valid correctness witnesses by using the novel technique for witness joining. To this end, we studied the following research questions:
@@ 7.1 Evaluation Setup
All experiments were run on machines with an Intel Xeon E3-1230 v5 @ 3.40 GHz (8 cores), GB of memory, and Ubuntu 22.04 LTS with Linux kernel 5.15.0. To increase the reproducibility of our experiments, we used BenchExec [30] for the execution. We used the same setup for each tool and instance of ranged program analysis: In a verification run, each configuration gets a task, containing a program and a specification, and is asked to either compute a proof (in case the program fulfills the specification) or to raise an alarm (if it violates the specification). We use the specification reach_error from the SV-Comp, which is fulfilled, if no assert in the program is violated111111In the dataset, a special function __VERIFIER_assert() is used to model assertions.. Each verification run is limited to a total of min CPU time on CPU cores and GB of memory, yielding a setup that is comparable to the one used in SV-Comp. We conducted the experiments on a subset of the SV-Benchmarks used in the SV-Comp and all experiments were conducted once. The subset contains in total C-tasks from all sub-categories of the SV-Comp category reach-safety [13], whereof tasks fulfill the property and the other contain a property violation. All data collected is available in our supplementary artifact [31].

### v2

@@ 4 Joining Witnesses
First, let us look at the verdict “false”, i.e., the ranged program analysis reports that the program violates the analyzed property. For the verdict “false”, it is sufficient that at least one of the ranged analyses reports “false”. This analysis also produces a violation witness [19, 20] that describes a counterexample path, which witnesses the violation. Furthermore, its violation witness is sufficient to explain the aggregated verdict “false” of the ranged program analysis. If more than one ranged program analysis computes a violation witness, reporting all paths violating the property may be beneficial for users and developers, as one path might be shorter or simpler than the other. We execute the ranged analyses in parallel and currently abort the other ranged analyses as soon as one ranged analysis detects a counterexample. Waiting for the other ranged analyses to (potentially) compute another counterexample requires defining and measuring such an additional timeout while ensuring that the ranged program analysis reports the known verdict “false” before it times out. To keep the concept of ranged program analysis simple, we decided against joining violation witnesses and simply using one of the violation witnesses provided by a ranged analysis with the verdict “false”. More concretely, we use the first violation witness being generated. Adding a configurable handling of multiple violation witnesses, e.g., by joining them or selecting one heuristically, is planned as future work.
@@
Next, let us consider the verdict “true”, i.e., the ranged program analysis proves the property. In this case, all ranged analyses successfully proved that their subset of the program behavior fulfills the property. To explain their results, they each provide a correctness witness [12, 20]. Formally, a correctness witness for a CFA is a (protocol) automaton , which uses the set of all quantifier-free predicate-logic formulae over the program variables to describe sets of data states. Similar to most automata, a correctness witness consists of a set of states including initial state , an alphabet whose symbols define which operations (i.e., CFA edges) and data states are considered by an automaton transition, and a transition relation . In addition, it contains a total function that assigns to each state of the automaton an invariant describing an expected set of data states. Furthermore, a correctness witness must fulfill the following completeness property, which guarantees that the correctness witness does not exclude any program behavior.
@@
Figure 4 shows three correctness witnesses for our example program. For this example, their graph structure is identical and reflects the CFA structure. However, the witnesses differ in their invariants. The invariants are the second part of the node labels and are shown in a different color for each witness. Note that we use o/w (short for otherwise) to summarize all control-flow edges that do not occur in another outgoing edge of the respective state. For example, for the initial state o/w stands for . The witnesses with the green and red invariants, respectively, both provide an explanation for the correctness of the subset of the program behavior. The witness with green invariants might be generated by a ranged analysis based on predicate abstraction [21] considering range [] resulting from test . The witness with the red invariants might be produced by a ranged analysis based on value analysis [22] considering the opposite range []. We observe that both witnesses use the invariant “false” for locations that the respective analysis never explored because no paths in their range contain these locations. Furthermore, note that the analyses use block encoding and, therefore, may only provide invariants for particular locations (either loop heads or locations with multiple outgoing edges). Locations, which are intermediate locations in the block encoding, are assigned the invariant true.
@@
Each of the two witnesses generated by a ranged analysis provides an explanation for the correctness of the subset of the program behavior analyzed by the respective ranged analysis but likely does not provide sufficient information for behavior analyzed by another ranged analyses. In the following, we explain how to combine correctness witnesses produced by ranged analyses during ranged program analysis into a joint correctness witness that explains the correctness of the complete program. The black invariants belong to such a joint witness. Next, we explain how to compute these joint witnesses.
Algorithm 1 describes the combination for two correctness witnesses999The algorithm can easily be extended to combine more than two witnesses. Alternatively, one can for example iteratively combine multiple witnesses by first joining the first two witnesses with Alg. 1 and then successively using Alg. 1 to combine the result of the first n-1 witness combinations with the n-th witness.. Its main part (lines 1–9) performs a parallel exploration of the program and the two witnesses.
@@
While a parallel exploration of the two witnesses would have been sufficient to compute a joint witness, we include the program into the exploration because it allows us to focus on explanations for program behavior, to construct (compact) joint witnesses that align with the program structure, which we assume are better for validation, and to simplify the witness construction. The parallel exploration starts in the initial location and initial automata states and explores the successors of each explored node once. Since we only need to provide explanations for program behavior, i.e., program paths, successor exploration focuses on successors that adhere to the program’s control-flow. To this end for each leaving control-flow edge in the program, we inspect all combinations of outgoing transitions of the witnesses that consider edge and that can be taken jointly, i.e., their enabling conditions do not exclude each other. The resulting successor combinations are then considered further during exploration. After the parallel exploration, the joint witness is computed. To avoid that the joint witness forces a validator to unroll the program, which may make it inefficient, Alg. 1 computes a rather natural representation of the joint correctness witness, which aligns with the program structure. Therefore, line 10 computes one joint invariant per program location. For each node detected during the parallel exploration, we combine the invariant information from both witnesses. We choose to combine the information via disjunction. Thus, we may lose information in case both witnesses consider the full behavior when computing their invariant information. However, disjunction ensures that we remain sound in case different behaviors are considered. Based on the combination, the joint invariant for a program location is derived from the disjunction of all nodes explored that refer to that location . We use a disjunction because different nodes likely represent different behaviors. Note that we include “false” in the disjunction to account for syntactically unreachable locations. After computing the invariants, line 11 determines the transitions of the joint witness. We add transitions to encode the program’s control-flow. Due to the format of a transition, we cannot add the control-flow edge directly, but add a transition that accepts the control-flow edge and that is always enabled. To make the witness complete, we add self-loops to each state101010More precisely, we add self-loops to each program location of the joint witness that cover all remaining edges, which do not adhere to the control-flow. Lastly, line 12 returns the joint witness, which reuses the program’s locations and initial location as states and initial states, using the same alphabet as the witnesses and the previously computed transitions and joint invariants.
@@
We implemented Alg. 1 for joining of witnesses presented above in the software analysis framework CPAchecker [13]. For the parallel exploration, we compose a CPA tracking the control-flow with one CPA per witness that tracks the transitions of that witness. To allow us to combine the witness information of the explored states, we wrap the above composition of CPAs into another CPA (called WrapperCPA), whose transfer relation forwards to the composition. We then use CPAchecker’s meta-reachability analysis instantiated with the above WrapperCPA to perform the parallel exploration. Thereafter, we apply CPAchecker’s witness generation on the result of the parallel exploration, which relies on the WrapperCPA to combine the witness information and computes the witness similar to Alg. 1.
@@
Finally, let us briefly discuss which invariants a ranged analysis should provide for locations it never explored. Basically, there are two options: “true” or “false”. In contrast to “false”, “true” is always a correct invariant with respect to the complete program. However, correctness witnesses of ranged analyses that only analyze a subset of the program behavior already do not guarantee correct invariants for the complete program for locations they visited during their analysis, especially if they do not analyze all paths to that location. They only encode information on behavior that they have seen. Thus, “false” would better fit to the information encoded at other seen locations. In addition, providing “true” would result in losing all correctness witness information provided by a different ranged analysis due to the disjunction. Hence, from the perspective of witness joining “false” is preferred. Since correctness witnesses were not designed for incomplete analysis and do not provide any requirements on how to handle unreached code, it depends on the verifier how it handles locations not analyzed. Hence, the quality and usefulness of a joint correctness witness depends on how the individual verifiers in the ranged program analysis handle non-analyzed locations and, therefore, may differ amongst verifiers.
@@ 6 Implementation
To show the advantages of ranged program analysis and to evaluate the two novel techniques of witness joining and work stealing, we implemented ranged program analysis and the two novel extensions within CoVeriTeam [24], a tool for specifying compositions of verification tools.
Ranged Analyses within CPAchecker. At first, we implement the composition of range reduction from Sec. 3.2 with an arbitrary configurable program analysis in the tool CPAchecker [13], making use of the existing composite pattern. More precisely, we build a novel range reduction analysis within CPAchecker, which combines the transfer relations of lower bound CPA and upper bound CPA as formalized in Sec. 3.2. Hence, it reuses elements from the value analysis, especially from the transfer relation, which are already implemented within CPAchecker. Our implementation is open-source and published in the CPAchecker repository. To ease reproducibility and practical applicability, we included the implementation also in our artifact archived at Zenodo [25].
Ranged Analyses not using CPAchecker.
For the use of non-CPA-based analyses, we refer to [26], where a ranged program analysis based on instrumentation is presented. It allows for using any off-the-shelf analysis. Instead of composing the analysis with a range reduction, a so-called range program is generated. For this, reduction [3] and program instrumentation are used.
@@
Ranged Program Analysis. The composition of ranged program analyses from Sec. 3 is implemented within the tool CoVeriTeam [24]. The tool allows building parallel and sequential compositions making use of existing program analyses, especially those from CPAchecker. To be able to use work stealing, i.e., to keep track of which program range is already successfully analyzed, we build a novel ranged program analysis component within CoVeriTeam. Its task is to (1) orchestrate the composition of ranged analyses, especially for work stealing, and (2) the aggregation of the partial results. The algorithm follows the description in Fig. 2 and Sec. 5 and receives as inputs the two ranged analyses Ranged Analysis 1 and Ranged Analysis 2, and a splitter. It first uses the splitter to generate the ranges. If the splitter fails, e.g., Lb3 cannot compute a test case, when the program does not contain a loop, we execute Ranged Analysis 1 on the interval . We extended CoVeriTeam by adding ranged analyses as a new type of verifier. Thus, it executes the two ranged analyses in parallel using the execution mechanism of CoVeriTeam. The algorithm collects the verdicts and witnesses computed by the ranged analyses and employs, if configured, work stealing. Moreover, the witnesses are joined. Therefore, the partial verification witnesses generated by Ranged Analysis 1 and Ranged Analysis 2 are collected first. For the witnesses join, we implemented Alg. 1 as a standalone component within CPAchecker, which is executed by CoVeriTeam. Our implementation is open-source and publicly available in the repository of CoVeriTeam as well as in the artifact [25].
Generating Ranges. We implemented the splitters Lb3 that is described in Sec. 3.4 as standalone components within CPAchecker. It generates test cases in the standardized XML-based Test-Comp test-case format111111https://gitlab.com/sosy-lab/test-comp/test-format/blob/testcomp23/doc/Format.md. Therefore, it builds the program execution tree demand-driven, meaning that only the path containing three loop unrollings is generated. Using the path formula, a representation of the path using predicates, an SMT-solver is called to find satisfiable assignments for the program input variables. These assignments are then used within the test case that is returned.
@@ 7 Evaluation
The evaluation of ranged symbolic execution conducted by Siddiqui and Khurshid [11] focuses on the performance increases gained for symbolic execution. We showed in [15] that ranged program analysis can increase the overall performance for finding property violations of symbolic execution by %. In addition, we compared the performance of the four different splitters introduced in Sec. 3.4, whereas Lb3 leads to the best overall performance. Moreover, the real overall execution time that is needed to solve large or complex tasks can be reduced by using ranged program analysis by up to 20%. When using ranged program analysis with BMC, it is observed that tasks that could not be solved using a single BMC instance can be solved when using BMC within ranged program analysis. We also evaluated one combination or ranged program analysis that uses two different analysis techniques, namely symbolic execution and predicate abstraction. We observed that the performance increases for symbolic execution, the one with the lower performance.
In the following, we are interested in a more systematic analysis of the effect of using ranged program analysis. The overall goal of the evaluation is to investigate, to which extent using ranged program analysis can increase the performance (i.e. effectiveness and efficiency) of analyses like predicate abstraction or symbolic execution. Hence, we compare ranged program analysis with the basic analysis employed. First, we want to investigate whether using two instances of the same analysis within ranged program analysis pays off. Thereafter, we study the effect of combining two conceptually different analyses and also analyze work stealing. Lastly, we study the question of whether our ranged program analysis is still capable of generating valid correctness witnesses by using the novel technique for witness joining. To this end, we studied the following research questions:
@@
7.1 Employed CPA-based Analysis.
Value Analysis. Value Analysis [22] (sometimes called constant propagation or explicit state model checking) tries to compute the concrete values for variables occurring in the program. For this, the exact value of each variable from a set (called precision) is tracked along each program path. Information for locations reachable via multiple program paths is not merged. The employed configuration does not track all variable values but uses counterexample-guided abstraction refinement (CEGAR) [27] to iteratively refine the precision until the program can either be proven safe or a violation is detected.
Predicate Abstraction. The predicate abstraction used is the standard one in CPAchecker, configured to compute an abstraction of the program using predicates with adjustable block encoding [21], abstracting only at loop heads. The predicates used for the analysis are determined by CEGAR [27], lazy refinement [29], and interpolation [30].
Bounded Model Checking. The configuration of bounded model checking (BMC) employed is iterative, meaning that the -th iteration inspects the behavior of the CFA by unrolling each loop to a bound of iterations. If no property violation can be found, is increased. For inspecting the behavior of the program, BMC encodes the unrolled CFA and the property as a boolean formula using the unified SMT-based approach for software verification [31]. Hereafter, the satisfiability of the formula that encodes the program up to the bound is checked to detect property violations. In case that is large enough that the full program is encoded, and no property violation is detected, the program is safe.
@@ 7.2 Evaluation Setup
All experiments were run on machines with an Intel Xeon E3-1230 v5 @ 3.40 GHz (8 cores), GB of memory, and Ubuntu 22.04 LTS with Linux kernel 5.15.0.
To increase the reproducibility of our experiments, we used BenchExec [32] for the execution.
We conducted the experiments on a subset of the SV-Benchmarks used in the SV-Comp for the property reachability and all experiments were conducted once.
The subset contains in total C-tasks from all sub-categories of the SV-Comp category ReachSafety and ConcurrencySafety dealing with the reachability of error labels [14], whereof tasks fulfill the property and the other contain a property violation.
In a so-called verification run a tool is given one the C-tasks, containing a program and a specification, and is asked to either compute a proof (in case the program fulfills the specification) or to raise an alarm (if it violates the specification).
We limit the available resources to a total of min CPU time on CPU cores and GB of memory, yielding a setup that is comparable to the one used in SV-Comp.
To ensure a comparison on equal ground, we use the same resource limitations for each tool.
This means that the basic analyses have the same resources available as the ranged program analyses.
As we employ two ranged analyses within a ranged program analysis, the available resources are split between the two ranged analyses.121212We evaluated configurations using three ranged analyses in parallel in [15].
All data collected is available in our supplementary artifact [25].